### v1

@@ 1 Introduction
The application area that we focus on in this paper is credit scoring — one of the oldest applications of data analytics. Lenders build credit scoring models to help adjudge the risk involved in granting a loan and decide on the terms of the loan and the interest rate (Thomas et al., 2017). Traditional credit scoring models use loan- or borrower-level data to assess a loan applicant’s default risk, thus treating borrowers as independent entities. While potential default correlation between borrowers has been acknowledged for some time, it is only more recently that this has started to be further investigated using network science. This is part of a broader trend in which credit scoring research increasingly focuses on improving the performance of existing credit scoring models through the incorporation of machine learning methods, and the inclusion of alternative data sources such as network data (Bravo & Óskarsdóttir, 2020).
@@
The structure of this paper is as follows. The next section discusses a selection of previous work on GNN, RNN, and credit risk modelling. Section 3 explains the methodology, multilayer networks, embeddings, and the models used in this paper. Section 4 sheds light on the data, dynamic networks, and the experiments in the paper. Section 5 presents the experimental results and highlights some discussion points relevant to the models. The final section summarizes our conclusions and suggests future work.
@@ 2.1 Graph neural networks
A second type of GNN, known as GAT, uses an attention mechanism to learn node-level representations (Veličković et al., 2018). The encoder-decoder-based neural machine translation system was outperformed by the attention mechanism in natural language processing (Bahdanau et al., 2015). Nowadays, attention models are widely utilized for document categorization (Pappas & Popescu-Belis, 2017), recommendation systems (Xiao et al., 2017), and the creation of image captions (Xu et al., 2015). The attention mechanism concentrates on a few selected relevant attributes while ignoring other irrelevant attributes (Bahdanau et al., 2015). The categories of attention models that are now in use are global and local attention (Luong et al., 2015), soft and hard attention (Bahdanau et al., 2015), and self-attention (Vaswani et al., 2017). Self-attention is an attention mechanism that is applied to compute a representation of a single sequence. In conjunction with RNN or convolutions, self-attention can be used in many applications like learning sentence representations (Lin et al., 2017) and machine reading (Cheng et al., 2016). The GAT is a type of GNN that applies an attention mechanism to graph-structured data, so as to classify nodes. It computes the hidden representation of each node by paying attention to its adjacent nodes and then applying a self-attention strategy. GAT achieved state-of-the-art results in both transductive (semi-supervised) and inductive (supervised) settings.
@@
Most of the GNN models have been proposed for static graph learning; however, over the past few years, several machine learning models capturing the structure and evolution of dynamic graphs have been introduced. A complete review of representation learning approaches for dynamic graphs is given in Kazemi et al. (2020), while a more specialized review of GNN-based approaches for dynamic graphs is provided by Skarding et al. (2021). There are many different approaches to modelling spatial and temporal information in graph-structured data. Diffusion Convolutional Recurrent Neural Network (DCRNN; Li et al., 2018) and Spatio-Temporal Graph Convolutional Networks (STGCN; Yu et al., 2018) analyse graph-structured data first and pass the results to sequence-to-sequence models or RNNs. Structural-RNN collects spatial and temporal data synchronously to associate the graph structure with temporal data so that it can apply RNNs on new graphs (Jain et al., 2016). Dynamic Graph Convolutional Networks (DGCN) propose a novel approach that combines RNNs and GCNs to learn long short-term dependencies together with the graph structure (Manessi et al., 2020). EvolveGCN is an approach for graph representation learning in dynamic graphs that uses an RNN to evolve the parameters of a GCN model. By doing so, EvolveGCN can capture the dynamics of the graph sequence without relying on node embeddings (Pareja et al., 2020). Temporal Graph Convolutional Network (T-GCN) is a novel method for real-time traffic forecasting that uses GCN to learn the complex topological structure of the urban road network for spatial dependence. It also employs RNN to capture the dynamic changes in traffic data for temporal dependence (Zhao et al., 2020). Temporal Graph Attention (TGAT) has been proposed for inductive representation learning on temporal graphs that uses a self-attention mechanism and a novel functional time encoding technique to efficiently aggregate temporal-topological neighbourhood features and learn time-feature interactions. TGAT can handle both node classification and link prediction tasks, and can be extended to include temporal edge features (Xu et al., 2020). JODIE has been proposed to predict future user-item interactions in domains such as e-commerce, social networking, and education. It is a coupled recurrent neural network model that learns the embedding trajectories of users and items through representation learning, and it introduces a novel projection operator to estimate the embedding of the user at any time in the future (Kumar et al., 2019). Dynamic Representation (DyRep) encodes evolving information over dynamic graphs into low-dimensional representations, namely as embeddings, using an inductive deep representation learning framework. The learned embeddings drive the dynamics of two fundamental processes: communication and association between nodes in dynamic graphs (Trivedi et al., 2019). Dynamic Self-Attention Network (DySAT) computes node representations by jointly using self-attention layers along two dimensions: structural neighbourhood and temporal dynamics, and has been evaluated on link prediction experiments (Sankar et al., 2019).
@@
All the aforementioned methods have been developed for networks that are either static, or dynamic, but single layer. A unified approach that can handle networks that are both dynamic and multilayer has not been put forward yet. This is significant, as most real-life networks share both characteristics simultaneously. Our approach focuses on solving this problem.
2.2 Recurrent neural networks
RNNs have been widely used for dynamic learning. The main characteristic of the RNN architecture is the use of a cyclic connection that makes it capable of updating any state based on past states and current input data. RNNs can process examples one at a time, retaining a state that reflects a long context window. Traditional approaches such as fully RNNs (Elman, 1990) and selective RNNs (Šter, 2013), consisting of standard recurrent cells, have shown acceptable performance in some settings, such as language translation (Sutskever et al., 2014), speech recognition (Graves et al., 2007), and human activity recognition (Donahue et al., 2017). However, these models are inefficient as one character is fed into and output from the model at a time. They also suffer from another drawback, i.e., that they can only consider a limited number of contexts. So, when the time gap between the relevant input data is large, traditional RNNs show poor performance due to vanishing/exploding gradient.
Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and Gated Recurrent Unit (GRU) (Cho et al., 2014) are two variants of RNN that were developed to solve these problems. LSTM can be trained to link time lags of more than 1000 discrete time steps by implementing constant error flow within special units (Hochreiter & Schmidhuber, 1997). LSTM has many successful applications, including time series prediction (Lindemann et al., 2021), speech recognition (Oruh et al., 2022), grammar learning (Zeng & Cao, 2023), and human action recognition (Muhammad et al., 2021). GRU is a newer version of RNN that uses a simpler way of storing information than LSTM, by using fewer gates (Cho et al., 2014). Both LSTM and GRU use gated mechanisms to memorize as much information as possible; however, there are some differences between these two models (Chung et al., 2014). Comparing these two, LSTM has a more complex structure, more parameters, and longer training time. LSTM is known to be able to deal with long-range dependencies, making it the preferred choice for models built over data of relatively large size (Yang et al., 2020). As our data is of medium size and the interface between GNNs and RNNs is not fully explored, the choice between LSTM and GRU is not obvious. We will, therefore, compare both models in Section 5.
@@ 3.2 Topological embedding
Capturing the topological dependence in a network is a key problem, as neighbouring nodes could influence each other. In this work, we trial both GCN (Kipf & Welling, 2017) and GAT (Veličković et al., 2018) to obtain the topological relationship between a node and its neighbours, encode the topological structure of the network and the features of nodes, capturing the information within the node connections. GCN or GAT is applied to each $G^{(t)}$ to obtain a hidden representation matrix $Z^{(t)}$. Each row of $Z^{(t)}$ contains a node embedding, meaning that for node $v_{i}$ we have a sequence of embeddings $[Z_{i}^{(1)},Z_{i}^{(2)},...,Z_{i}^{(\tau)}]$.
@@ 3.3 Temporal embedding
Dealing with the temporal dependence is another key problem, as the temporal sequence of connections between nodes could provide useful information. As explained before, in this work, we use LSTM (Hochreiter & Schmidhuber, 1997) and GRU (Cho et al., 2014) to catch the information related to the evolution of the networks. After obtaining the sequence of topological embeddings $[Z_{i}^{(1)},Z_{i}^{(2)},...,Z_{i}^{(t)}]$, we need to input it into the RNN model and use the hidden representation of the RNN model as the temporal node embeddings for $v_{i}$.
LSTM uses a total of three gates, i.e., input gate, forget gate, and output gate. The input gate determines what information from the current topological embedding and previous temporal embeddings will be cached in long term memory. The forget gate decides which information from the long term memory should be maintained or repudiated. The output gate takes the current topological embedding, the previous temporal embedding and the newly computed long term memory to produce the new temporal embedding that will be passed on to the cell in the next time step. The LSTM model can be formulated as follows:
@@ 3.5 GNN-RNN-ATT models
The GNN-RNN-ATT model is another proposed model for DYMGNN. For GNN-RNN-ATT models, a soft attention mechanism is applied to assign different importance to each timestamp. Our approach for creating a new hidden state for the node embeddings that is more expressive of the global variation trends can be formulated as follows:
@@ 3.6 Decoder and loss function
A deep neural network model is typically comprised of an encoder and a decoder. The encoder takes input and produces embeddings, whereas the decoder takes the embeddings and performs the prediction task. In our specific case, GNN, RNN, and ATT comprise the encoder of the full model, while the decoder is a set of feed-forward neural networks applied to the node embeddings, followed by a series of layers that either apply a chosen activation function for non-linearity or dropout function for regularization. The final output is the model prediction for our binary outcome (here, default Y/N), i.e., whether the node $v_{i}$ belongs to class $1$ ($Y_{i}=1$) or 00 ($Y_{i}=0$). The decoder outputs a vector $\hat{Y}$ where $\hat{Y}_{i}$ specifies the probability of a node $v_{i}$ belonging to class $1$ given the snapshots $[G^{(1)},...,G^{(\tau)}]$. While there is no unique format for the decoder, the architecture used in this work is shown in Fig. 5.
@@
One of the most important aspects of a deep learning model is its loss function. For our work, we use the well-known binary cross-entropy loss function which can be written as
@@ 4.2 Dynamic networks
As we are interested in studying the effect of connections between the borrowers and the evolution of those connections over time, we use the data to create a sequence of dynamic networks following the process in subsection 3.1. In particular, we are interested in predicting one-year-ahead loan default based on application information and six months of borrower’s repayment behaviour.
For this work, loans originated in 2009 and 2010 are used for training and testing. This is to give loans sufficient time to default. We use application data and 18 months of behavioural data, from January 2012 to June 2013, for training. We also use application data and six months of behavioural data of a holdout set, from July 2013 to December 2013, for testing. We consider rolling windows, shifting by one month, for training and testing, with each window containing six snapshots $[G^{(1)},...,G^{(6)}]$, and each snapshot corresponding to one month. So, we have 13 windows for training and one window for testing. All snapshots of a specific window have the same set of nodes. However, the node set could be different from one window to another; a loan that has defaulted will remain marked as a defaulter for the observation window but will disappear once the window moves past it. During the training of each window, the goal is to predict default within 12 months following the month of the last snapshot in that window. A one-year horizon is practical for credit management and decision-making, as it balances the need for a sufficiently long period to assess risk while not extending so far that predictions become highly speculative (Lopez & Saidenberg, 2000). Fig. 6 displays the timeline for the windows and their corresponding horizons in model training.
@@
We can train the models using either single layer or double layer networks, with geographical location of the borrower and the company lending the loan being the connector variables. Borrowers whose postcodes have the same first two digits are assumed to be in the same geographical area. In the case of the double layer network, nodes in one layer are also connected to their twins in the other layer. Some previous studies showed that applying some sort of dropout techniques on graph structures could help increase the expressiveness of the GNN models (Shu et al., 2022). Therefore, we decide to randomly select and isolate 50% of the nodes in each snapshot of a window. In other words, in each snapshot, at least half of the nodes do not have any connections with other nodes.
@@
Training
148,520
16,368,244
148,520
91,486,176
297,040
108,151,460
Validation
82,180
4,725,842
82,180
27,735,664
164,360
32,625,866
Test
96,490
6,761,051
96,490
38,404,277
192,980
45,358,308
@@ 5.1 Baseline methods
We benchmark our proposed model against a selection of GNN-based and non-GNN-based baseline models. Table 3 shows the results for two GNN-based baseline models, i.e., static GCN and static GAT, whereas Table 4 shows the results for three non-GNN-based baseline models, i.e., Logistic Regression (LR), XGBoost (XGB), and a Deep Neural Network (DNN). LR is popular in the commercial and financial sectors due to its straightforwardness and ease of understanding. Meanwhile, XGB has established itself as a powerful technique for both classification and regression tasks involving structured datasets (Lessmann et al., 2015). DNN is fundamental to deep learning and has seen broad usage across a variety of predictive tasks.

### v2

@@ 1 Introduction
The application area that we focus on in this paper is credit scoring — one of the prominent applications of data analytics. Lenders build credit scoring models to help adjudge the risk involved in granting a loan and decide on the terms of the loan and the interest rate (Thomas et al., 2017). Traditional credit scoring models use loan- or borrower-level data to assess a loan applicant’s default risk, thus treating borrowers as independent entities. While potential default correlation between borrowers has been acknowledged for some time, it is only more recently that this has started to be further investigated using network science. This is part of a broader trend in which credit scoring research increasingly focuses on improving the performance of existing credit scoring models through the incorporation of machine learning methods, and the inclusion of alternative data sources such as network data (Bravo & Óskarsdóttir, 2020).
@@
The structure of this paper is as follows. The next section discusses a selection of previous work on GNN and credit risk modelling. Section 3 explains the methodology, multilayer networks, embeddings, and the models used in this paper. Section 4 sheds light on the data, dynamic networks, and the experiments in the paper. Section 5 presents the experimental results and highlights some discussion points relevant to the models. The final section summarizes our conclusions and suggests future work.
@@ 2.1 Graph neural networks
A second type of GNN, known as GAT, uses an attention mechanism to learn node-level representations (Veličković et al., 2018). The encoder-decoder-based neural machine translation system was outperformed by the attention mechanism in natural language processing (Bahdanau et al., 2015). Nowadays, attention models are widely utilized for document categorization (Pappas & Popescu-Belis, 2017), recommendation systems (Xiao et al., 2017), and the creation of image captions (Xu et al., 2015). The attention mechanism concentrates on a few selected relevant attributes while ignoring other irrelevant attributes (Bahdanau et al., 2015). The categories of attention models that are now in use are global and local attention (Luong et al., 2015), soft and hard attention (Bahdanau et al., 2015), and self-attention (Vaswani et al., 2017). Global attention considers all source positions when deriving the context vector, providing a comprehensive overview of the entire input sequence. In contrast, local attention focuses on a subset of source positions, making it more efficient by narrowing the context to relevant parts. Soft attention generates a weighted sum of the attention scores, allowing gradients to pass through and making the model end-to-end trainable. Hard attention, on the other hand, selects a single source position, making it less computationally expensive but more challenging to optimize due to its non-differentiable nature. Self-attention is an attention mechanism that is applied to compute a representation of a single sequence, enabling the model to weigh the importance of different positions in the sequence when generating its output. In conjunction with RNN or convolutions, self-attention can be used in many applications like learning sentence representations (Lin et al., 2017) and machine reading (Cheng et al., 2016). The GAT is a type of GNN that applies an attention mechanism to graph-structured data, so as to classify nodes. It computes the hidden representation of each node by paying attention to its adjacent nodes and then applying a self-attention strategy. GAT achieved state-of-the-art results in both transductive (semi-supervised) and inductive (supervised) settings.
@@
Most of the GNN models have been proposed for static graph learning; however, over the past few years, several machine learning models capturing the structure and evolution of dynamic graphs have been introduced. A complete review of representation learning approaches for dynamic graphs is given in Kazemi et al. (2020), while a more specialized review of GNN-based approaches for dynamic graphs is provided by Skarding et al. (2021). There are many different approaches to modelling spatial and temporal information in graph-structured data. Diffusion Convolutional Recurrent Neural Network (DCRNN; Li et al., 2018) and Spatio-Temporal Graph Convolutional Networks (STGCN; Yu et al., 2018) analyse graph-structured data first and pass the results to sequence-to-sequence models or RNNs. Structural-RNN collects spatial and temporal data synchronously to associate the graph structure with temporal data so that it can apply RNNs on new graphs (Jain et al., 2016). Dynamic Graph Convolutional Networks (DGCN) propose a novel approach that combines RNNs and GCNs to learn long short-term dependencies together with the graph structure (Manessi et al., 2020). EvolveGCN is an approach for graph representation learning in dynamic graphs that uses an RNN to evolve the parameters of a GCN model. By doing so, EvolveGCN can capture the dynamics of the graph sequence without relying on node embeddings (Pareja et al., 2020). Temporal Graph Convolutional Network (T-GCN) is a novel method for real-time traffic forecasting that uses GCN to learn the complex topological structure of the urban road network for spatial dependence. It also employs RNN to capture the dynamic changes in traffic data for temporal dependence (Zhao et al., 2020). Temporal Graph Attention (TGAT) has been proposed for inductive representation learning on temporal graphs that uses a self-attention mechanism and a novel functional time encoding technique to efficiently aggregate temporal-topological neighbourhood features and learn time-feature interactions. TGAT can handle both node classification and link prediction tasks, and can be extended to include temporal edge features (Xu et al., 2020). Joint Dynamic User-Item Embeddings (JODIE) has been proposed to predict future user-item interactions in domains such as e-commerce, social networking, and education. It is a coupled recurrent neural network model that learns the embedding trajectories of users and items through representation learning, and it introduces a novel projection operator to estimate the embedding of the user at any time in the future (Kumar et al., 2019). Dynamic Representation (DyRep) encodes evolving information over dynamic graphs into low-dimensional representations, namely as embeddings, using an inductive deep representation learning framework. The learned embeddings drive the dynamics of two fundamental processes: communication and association between nodes in dynamic graphs (Trivedi et al., 2019). Dynamic Self-Attention Network (DySAT) computes node representations by jointly using self-attention layers along two dimensions: structural neighbourhood and temporal dynamics, and has been evaluated on link prediction experiments (Sankar et al., 2019).
@@
All the aforementioned methods have been developed for networks that are either static and single layer, static and multilayer, or dynamic and single layer. A unified approach that can handle networks that are both dynamic and multilayer has not been put forward yet. This is significant, as most real-life networks share both characteristics simultaneously. Our approach focuses on solving this problem.
@@ 3.2 Topological embedding
Neural networks are a type of machine learning model inspired by the human brain, consisting of layers of interconnected neurons. Each neuron processes input data, applies a mathematical transformation using weights and biases, and passes the result to the next layer. They are trained through a process called backpropagation, which adjusts the weights and biases to minimize prediction errors measured by a loss function (Goodfellow et al., 2016). GNNs extend these neural networks to handle graph-structured data.
Capturing the topological dependence in a network is a key problem, as neighbouring nodes could influence each other. In this work, we trial two different types of GNNs, i.e., GCN (Kipf & Welling, 2017) and GAT (Veličković et al., 2018), to obtain the topological relationship between a node and its neighbours, encode the topological structure of the network and the features of nodes, capturing the information within the node connections. GCN or GAT is applied to each $G^{(t)}$ to obtain a hidden representation matrix $Z^{(t)}$. Each row of $Z^{(t)}$ contains a node embedding, meaning that for node $v_{i}$ we have a sequence of embeddings $[Z_{i}^{(1)},Z_{i}^{(2)},...,Z_{i}^{(\tau)}]$.
@@ 3.3 Temporal embedding
Dealing with the temporal dependence is another key problem, as the temporal sequence of connections between nodes could provide useful information. In this work, we use LSTM (Hochreiter & Schmidhuber, 1997) and GRU (Cho et al., 2014) to capture the information related to the evolution of the networks. Both LSTM and GRU use gated mechanisms to memorize as much information as possible; however, there are some differences between these two models (Chung et al., 2014). Comparing these two, LSTM has a more complex structure, more parameters, and longer training time. LSTM is known to be able to deal with long-range dependencies, making it the preferred choice for models built over data of relatively large size (Yang et al., 2020). As our data is of medium size and the interface between GNNs and RNNs is not fully explored, the choice between LSTM and GRU is not obvious. We will, therefore, compare both models in Section 5. After obtaining the sequence of topological embeddings $[Z_{i}^{(1)},Z_{i}^{(2)},...,Z_{i}^{(t)}]$, we need to input it into the RNN model and use the hidden representation of the RNN model as the temporal node embeddings for $v_{i}$.
LSTM uses a total of three gates, i.e., input gate, forget gate, and output gate. The input gate determines what information from the current topological embedding and previous temporal embeddings will be cached, or stored for future use, in long term memory. The forget gate decides which information from the long term memory should be maintained or repudiated. The output gate takes the current topological embedding, the previous temporal embedding and the newly computed long term memory to produce the new temporal embedding that will be passed on to the cell in the next time step. The LSTM model can be formulated as follows:
@@ 3.5 GNN-RNN-ATT models
The GNN-RNN-ATT model is another proposed model for DYMGNN. In GNN-RNN-ATT models, a soft attention mechanism is applied to assign different importance to each timestamp. This approach contrasts with GNN-RNN models which assign equal importance to every timestamp. The use of attention in GNN-RNN-ATT models allows for a more nuanced weighting of temporal information. Our approach for creating a new hidden state for the node embeddings that is more expressive of the global variation trends can be formulated as follows:
@@ 3.6 Decoder and loss function
A deep neural network model is typically comprised of an encoder and a decoder. The encoder takes input and produces embeddings, whereas the decoder takes the embeddings and performs the prediction task (Goodfellow et al., 2016). In our specific case, GNN, RNN, and ATT comprise the encoder of the full model, while the decoder is a set of feed-forward neural networks applied to the node embeddings, followed by a series of layers that either apply a chosen activation function for non-linearity or dropout function for regularization. The final output is the model prediction for our binary outcome (here, default Y/N), i.e., whether the node $v_{i}$ belongs to class $1$ ($Y_{i}=1$) or 00 ($Y_{i}=0$). The decoder outputs a vector $\hat{Y}$ where $\hat{Y}_{i}$ specifies the probability of a node $v_{i}$ belonging to class $1$ given the snapshots $[G^{(1)},...,G^{(\tau)}]$. While there is no unique format for the decoder, the architecture used in this work is shown in Fig. 5.
@@
One of the most important aspects of a deep learning model is its loss function. For our work, we use the well-known binary cross-entropy loss function (Gneiting & Raftery, 2007) which can be written as
@@ 4.2 Dynamic networks
As we are interested in studying the effect of connections between the borrowers and the evolution of those connections over time, we use the data to create a sequence of dynamic networks following the process in subsection 3.1. In particular, we are interested in predicting one-year-ahead loan default based on application information and six months of borrower’s repayment behaviour. We choose a six-month period because six and twelve months are common choices for lookback periods (Kennedy et al., 2013). Furthermore, this paper will later show that extending beyond six months offers minimal additional benefits.
For this work, loans originated in 2009 and 2010 are used for training and testing. We select these years to ensure sufficient default information is available for reliably comparing the models. It is important to note, however, that loan population and behaviour change over time, and our sample data may reflect some effects of the global financial crisis. Thus, further research could explore the robustness of our proposed methods across different time periods and financial conditions. We use application data and 18 months of behavioural data, from January 2012 to June 2013, for training. We also use application data and six months of behavioural data of a holdout set, from July 2013 to December 2013, for testing. We consider rolling windows, shifting by one month, for training and testing, with each window containing six snapshots $[G^{(1)},...,G^{(6)}]$, and each snapshot corresponding to one month. So, we have 13 windows for training and one window for testing. All snapshots of a specific window have the same set of nodes. However, the node set could be different from one window to another; a loan that has defaulted will remain marked as a defaulter for the observation window but will disappear once the window moves past it. During the training of each window, the goal is to predict default within 12 months following the month of the last snapshot in that window. A one-year horizon is practical for credit management and decision-making, as it balances the need for a sufficiently long period to assess risk while not extending so far that predictions become highly speculative (Lopez & Saidenberg, 2000). Fig. 6 displays the timeline for the windows and their corresponding horizons in model training.
@@
We can train the models using either single layer or double layer networks, with geographical location of the borrower and the company lending the loan being the connector variables. Borrowers whose zip codes have the same first two digits are assumed to be in the same geographical area. In the case of the double layer network, nodes in one layer are also connected to their twins in the other layer. Some previous studies showed that applying some sort of dropout techniques on graph structures could help increase the expressiveness of the GNN models (Shu et al., 2022). Therefore, we decide to randomly select and isolate 50% of the nodes in each snapshot of a window. In other words, in each snapshot, at least half of the nodes do not have any connections with other nodes.
@@
Training 148,520 16,368,244 148,520 91,486,176 297,040 108,151,460
Validation 82,180 4,725,842 82,180 27,735,664 164,360 32,625,866
Test 96,490 6,761,051 96,490 38,404,277 192,980 45,358,308
@@ 5.1 Baseline methods
We benchmark our proposed model against a selection of GNN-based and non-GNN-based baseline models. Table 3 shows the results for two GNN-based baseline models, i.e., static GCN and static GAT, whereas Table 4 shows the results for three non-GNN-based baseline models, i.e., Logistic Regression (LR), XGBoost (XGB), and a Deep Neural Network (DNN). LR is popular in the commercial and financial sectors due to its straightforwardness and ease of understanding. Meanwhile, XGB has established itself as a powerful technique for both classification and regression tasks involving structured datasets (Gunnarsson et al., 2021). DNN is fundamental to deep learning and has seen broad usage across a variety of predictive tasks.